OpenStudy (anonymous):

20-(4/5)x is greater than or equal to 16. How do I solve?

OpenStudy (anonymous):

\[20-(4/5)x \ge16\] step 1: subtract 16 both sides & add (4/5)x to both sides \[20-16\ge(4/5)x\] \[4\ge(4/5)x\] step 2:multiply/divide to get rid of the coefficient, 4/5 \[5/4 *4\ge x\]\[5\ge x\]